OpenStudy (anonymous):

Let y = 1/x^4. Find values A and B so that (dy)/(dx) = A/(x^B). A = B =

myininaya (myininaya):

\[y=x^{-4} => y'=(-4)x^{-4-1}=-4 \cdot \frac{1}{x^5}=\frac{-4}{x^5}\]

myininaya (myininaya):

can you identify A and B now?

OpenStudy (anonymous):

is the formula to solve for a?

OpenStudy (anonymous):

what do I need to do after -4/ x^5?

myininaya (myininaya):

identify A and B

myininaya (myininaya):

\[y'=\frac{-4}{x^5}=\frac{A}{x^B}\]

myininaya (myininaya):

so A=? and B=?

OpenStudy (anonymous):

-4=a? and

OpenStudy (anonymous):

b= x^5?

myininaya (myininaya):

A=-4 is right but B does not equal x^5

OpenStudy (anonymous):

b= 5

myininaya (myininaya):

B=5